OpenStudy (anonymous):

what is the derivative of f(x)=-(x-1)^2-2 WITHOUT using the chain rule only the product rule or the quotient rule?

OpenStudy (amistre64):

why not just use first principles then and ignore all the rules :)

OpenStudy (amistre64):

-(x-1)^2-2 -(x^2-2x+1)-2 -x^2 + 2x - 3 now its only power and constant rules ....

OpenStudy (anonymous):

Using the product rule: \[f(x)=-(x-1)(x-1)-2\] \[f'(x)=-(x-1)(x-1)'+(-(x-1))'(x-1)+(2)'\]

OpenStudy (anonymous):

' is prime so take the derivative where you need to and simplify recall that (2)'=0 because the derivative of any constant(number) is 0 with respect to x (or any other variable for that matter).

OpenStudy (anonymous):

well what i have is... f(x)=-(x-1)^2-2 f(x)=-(x-1)-2 f(x)=-x+1-2 But then by my instructors example she has this random example similar to mine... and so then she obviously has the derivative of f(x) but Spring break just ended and I totally feel retarded because I dont remember how to get the derivative of the f(x) we havent gone over the chain rule so when I look for examples many use the chain rule so IM stuck

OpenStudy (phi):

you said ********** well what i have is... f(x)=-(x-1)^2-2 f(x)=-(x-1)-2 f(x)=-x+1-2 **************** I don't know what that means. Is this 3 different problems ?

OpenStudy (anonymous):

No Phi I believe i answered the problem wrong. the answer from f(x)=-(x-1)^2-2 should be f(x)=x^2+2x-3... right?

OpenStudy (phi):

you can expand (x-1)(x-1) to get x^2 -2x +1 (use FOIL) you can put that into f(x)= - (x-1)^2-2 don't lose track of the minus sign out front. you get f(x)= - ( x^2 -2x +1 ) -2 you can distribute the -1 to get f(x)= -x^2 +2x-1 -2 and combine -1-2 so we get f(x)= -x^2 +2x-3

OpenStudy (phi):

But I would learn to do it this way f(x)=-(x-1)^2-2 let u= (x-1) then the problem is \[ f(u) = -u^2 -2 \] now take the derivative term by term \[ d f(u)= - d\ u^2 - d\ 2 \] the derivative of the constant 2 is zero so we only have to worry about the first term d u^2 = 2 u du and \[ d f(u)= -2 u du\] remember u= x-1 du = d( x-1) term by term: du = dx - 0= dx we get (replacing u with (x-1) \[d(\ -(x-1)^2-2) = -2 (x-1) dx \] and \[ \frac{d}{dx} f(x)= -2 (x-1) \]